OpenStudy (anonymous):

In a small bag of candy, there are 10 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(You select first an orange candy and then a blue candy).

OpenStudy (lgbasallote):

probability you'll get an orange candy first is 10/14 probability you'll get a blue candy next is 4/13 (because you got i candy already) so multiply \[\large \frac{10}{14} \times \frac{4}{13}\]

OpenStudy (anonymous):

how do you multiply the fractions?

OpenStudy (lgbasallote):

multiply the numerator then multiply the denominator that gives you \[\large \frac{10 \times 4}{14 \times 13}\] something can be canceled right?

OpenStudy (anonymous):

i got it, thanks.

OpenStudy (anonymous):

20/91

OpenStudy (lgbasallote):

seems right